In soybeans, dwarf (D) is dominant to tall (d), round (R) is dominant to oval (r), and green (G) is dominant to brown (g). The genes are linked and the map is given below. Pure-breeding tall, round, green plants were mated to pure-breeding dwarf, oval, brown plants. The F1 were testcrossed. All loci are autosomal. The coefficient of coincidence is 1.

Height Shape Color
14cM 22cM

Rank the progeny phenotypes below to indicate their frequency in the testcross progeny based on the data above. Use 1 to indicate the most frequent then 2, then 3 and let 4 indicate the least frequent.

tall, oval, green

tall, oval, brown

tall, round, brown

tall, round, green
Answer #1

Answer:

Tall, oval, green----4 = 1.5%

Tall, oval, brown---3 = 5.5%

Tall, round, brown—2 = 9.5%

Tall, round, green---1 = 33.5%

Explanation:

dRG/dRG x DrG/Drg –Parents

dRG/Drg---------------F1

dRG/Drg x drg/drg ---Testcross

Explanation:

d---14cM---r----22cM--g

Distance between the genes (mu) = (%) of recombination frequency

Expected double crossover frequency = 22% x 14% = 0.0308

Coefficient of coincidence = obsereved dco / expected dco

1 =?/0.0308

Observed dco = 1 * 0.0308 = 0.0308 = 3%

Explanation:

Class of gametes

Phenotype

Frequency of reciprocal pair

Numbers

Frequency of each class

Phenotypes

Parental

Drg

1- all recombinant progeny

100%-(19%+11%+3%) = 67%

33.5%

Dwarf, oval, brown

dRG

33.5%

Tall, round, green

SCO 1

DRG

RF in region 1 = SCO1 - DCO

SCO1 = 14%-3% = 11%

5.5%

Dwarf, round, green

Drg

5.5%

Tall, oval, brown

SCO 2

DrG

RF in region 2 = SCO2 - DCO

SCO2 = 22% – 3% = 19%

9.5%

Dwarf, oval, green

dRg

9.5%

Tall, round, brown

DCO

DRg

(RF in region 1) * (RF in region 2)

3%

1.5%

Dwarf, round, brown

DrG

1.5%

Tall, oval, green